# A proof-of-principle experiment on the spontaneous symmetry breaking machine and numerical estimation of its performance on the benchmark problem

### v1

@@ Abstract
In a previous paper, we proposed a unique physically implemented type simulator for combinatorial optimization problems, called a spontaneous symmetry breaking machine (SSBM). In this paper, we first report the results of the experimental verification of SSBM using a small-scale benchmark system, and then describe numerical simulations using the benchmark problems () conducted to confirm its usefulness for large-scale problems. From 1000 samples with different initial fluctuations, it became clear that SSBM can explore a single extremely stable state. This is based on the principle of a phenomenon used in SSBM, and could be a notable advantage over other simulators.
@@ I Introduction
Symmetry and its (spontaneous) breaking is one of the most significant concepts for understanding various natural phenomena, and it is widely known that they play an essential role in a wide range of areas covered by science [1]. In applied physics, which is at the heart of scientific applications, and in condensed matter physics, which is the main foundation of applied physics, there are many arguments suggesting that spontaneous symmetry breaking (SSB) may play an important role in various phenomena. In those, the discussion is mainly focused on explaining phase transitions (e.g. spontaneous magnetization [2], insulator-metal phases [3]) and changes of state before and after phase transitions, and the main emphasis is on understanding the manifestation of macroscopic properties based on the assumption of the existence of microscopic elements (e.g. electron spin). In dissipative systems, concepts such as dissipative structure and synergetics have been discovered, leading to a broader and deeper understanding of natural phenomena from the perspective of dynamical macroscopic orders created by symmetry breaking [4, 5]. In the field of elementary particle theory, it goes even further and plays an important role in understanding the manifestation of elementary particles (microscopic elements) themselves such as the Higgs boson [6, 7].
In recent years, there has been growing interest in attempts to solve problems in an extremely wide range of fields (logistics, financial portfolios, drug discovery, machine learning, etc.) by treating them as combinatorial optimization problems (COPs). At the same time, there has been a lot of research on solving Ising problems that can be mapped to COPs in the NP-hard category [8]. In particular, quantum annealers (a type of quantum computer) [9, 10] and coherent Ising machines (based on laser oscillation phenomena) [11] that use physically implemented type simulators to explore the ground state of Ising model systems (IMS) are attracting attention.
@@ II SSB machine
Polarized coherent clock pulses with pulse width and repetition interval are input to the optical input port of a 12 MZM used as a partial inflow gate. The -th output light pulse from output port A of the MZM is converted into an electrical pulse by a photodiode (PD) after passing through an optical variable delay line to adjust the delay and timing. This electrical pulse is then expanded to a pulse width of using a Bessel filter, while its power is compensated by two amplifiers while suppressing excessive noise, and is used for intensity modulation by the MZM for the ()-th optical input pulse. In the above operation of this circuit, we can identify independent fully dissipative systems defined by virtual boundaries where the -th and ()-th clock pulses interact, with the MZM serving as the common inflow gate. And, when the condition is satisfied, this fully dissipative system maintains an extremely singular state described by the following iterative equation using the parameters phase difference-induced efficiency , the transmittance and to port A at the input of the clock pulses and ()-th, and the static phase condition (offset) , related to MZM, thus establishing the dissipative causality.
Here, we set the remaining degrees of freedom as and . These are the parameters that enable the SSB phenomenon to occur in the SSBM [12]. However, when the optical clock pulse is supplied and the dissipative causality is established, the states appear and are maintained (becoming hidden states). Therefore, it is necessary to input an initialization optical pulse train from the optical input port with adjusted timing to transition all full-dissipative systems to an unstable fixed point (), which is the state before SSB begins (SBSB). Moreover, we will consider the construction of a many-body-like system by imparting interrelationships between full-dissipative systems capable of causing the SSB phenomenon described above. Specifically, we treat the two attractors in the above full-dissipative systems as pseudo spin up and spin down, and introduce pseudo-spin interactions (PSIs) represented by the following equations, which can be implemented by an optical interference circuit.
@@
As is clear from the equations these PSIs are different from the so-called exchange interactions, but they satisfy the behavior consistent with the interpretation of exchange energy (namely, ferromagnetic types stabilize when states are aligned, while antiferromagnetic types stabilize when states are opposite). This is necessary to deal with the fact that pseudo-spins are continuous variables rather than discrete variables. Even though pseudo-spins are continuous variables, if we use interactions that follow exchange interactions as they are, we cannot guaranty the expected behavior in regions other than those corresponding to the values assigned to spins. The introduction of these PSIs yields new and interesting results. As is clear from Equations (4) and (5), when full-dissipative systems are simultaneously in the SBSB , the effects of PSIs disappear. Therefore, it is possible to transition the entire complex system consisting of full-dissipative systems with PSIs to SBSB through the initialization process described above. Moreover, this complex system creates a many-body-like system consisting of pseudo-spins, incorporating the effects of PSIs through the SSB phenomenon. We refer to a many-body-like system composed of pseudo-spins as described above as an Ising model like system (IMLS).
@@ III Proof-of-principle experiment
To understand the creation of the IMLS through the SSB phenomenon, we cannot simply apply the concepts developed in condensed matter physics. And we do not know of any other specific examples of the phenomenon of many-body-like systems being created by SSB. However, since consistency with the interpretation of exchange interaction energy was ensured when introducing PSIs, it is expected that a correspondence will exist between the stable states in the IMS and those in the IMLS. In our previous paper, we conducted numerical simulations to confirm this correspondence and found it promising [12]. Here, we present one of the most compelling pieces of evidence to ensure the reliability of this correspondence (duality) by experimentally demonstrating the effectiveness of SSBM as a simulator for COPs.
@@
Figure 3 is a third-order graph representing a COP belonging to the category called MaxCut3 [15], which is addressed in the proof-of-principle experiment. This COP is an established problem that was used in demonstration experiments of coherent Ising machines, which are one type of physically implemented Ising machine. And this COP has the advantage that dynamic switching is not required in the optical delay interference circuit (ODIC) [23] for the physical implementation of pseudo-spin interactions.
@@
Figure 4(a) shows a schematic diagram of the ODIC, and Fig.4(b) shows a schematic diagram of a proof-of-principle system for SSBM that implements this circuit. The ODIC has input ports corresponding to the output port and the complementary output port of the 12 MZM, and the optical delay interference between the input optical pulses from these input ports physically implements the relationships (pseudo-spin interactions) between the 16 full-dissipative systems, which is then output to the variable optical delay line in the subsequent stage. This ODIC is designed so that, under the condition that no additional optical amplification is performed, the losses for branching and adjustment (phase, attenuation) are taken into account to satisfy . The delay of the third route corresponding to is set to be the smallest, and the delays of the routes corresponding to the reference route (main route and second route:), and are set to be times larger than the clock pulse interval. In addition, by adjusting the phase shifters and attenuators, the pseudo-spin interaction represented by Eqns.(2)-(7) is physically implemented through optical delay interference. A series of optical clock pulses with a repetition frequency of 1.02 GHz is extracted from a series of optical output pulses from an active mode-locked laser with a repetition frequency of 8.16 GHz and a pulse width of 10 ps and then split into two. To allow repeated trials, OCPs with a width of and a period of 0.996 MHz are extracted from one of these optical pulse trains and reset regions with a width of are set. From the other of these optical pulse trains, a sequence consisting of the first 16 consecutive optical pulses is extracted and prepared as an optical initializing pulse train (OIPT). In addition, we adjusted the timing so that the electrical pulse originating from the first optical pulse in this OIPT is input as the intensity-modulated signal to the 12 MZM at the same time as the first optical pulse of the OCPs is input to the 12 MZM. With the above settings, SSBM will reset and immediately transition to the NS, causing the SSB phenomenon to start, repeating this behavior. The peak power of the inflow optical pulse detected by the PD corresponds to the transmittance of each system, so this electrical pulse was normalized to monitor the behavior of the system. Using a real-time oscilloscope, a series of 800 trials was recorded at once.
@@
The waveforms shown in Fig. 5(a) correspond to the case where the PSI is turned off (set ) using attenuators in the ODIC. It was confirmed that the fundamental SSB phenomenon occurs in each of the 16 independent full-dissipative systems, where the system transitions to either the 0-state or 1-state in a manner that reflects the initial fluctuations. In cases without PSI, since full-dissipative systems are entirely independent and do not compete with each other, state transitions are completed in approximately 20 steps, with little variation in the number of steps required. Figures 5(b)-(c) show the normalized monitoring signals of the SSBM with pseudo-spin interactions. In this example, it can be confirmed that the state transition is completed in approximately 30 steps, successfully transitioning to the state corresponding to the most stable solution. Furthermore, the number of steps required for solution search exhibited a relatively large variation and required significantly more steps compared to when PSI was off, with some cases requiring over 50 steps. This variability is believed to stem from the fact that PSI-induced conflicts exhibit different behaviors in individual cases. Based on data obtained from experiments, when thresholding was performed defining values above 80% of the normalized value as up and values below 20% as down, the success rate to find the most stable solution was approximately 97%. These results demonstrate that SSBM is a promising physical implementation simulator for combinatorial optimization problems. They also strongly support the existence of a duality (correspondence) between the IMLS states that appear in SSBM and the stable states of IMS.
@@ IV.1 Asymmetric Effects in Pseudo-Spin Interactions
Figures 6 and 7 show the results of numerical simulations for the MaxCut3 problem and a fully connected complete graph problem, respectively. The significant difference in the magnitude of the PSI employed for these two problems is due to the differing densities of coupling, and the calculations were carried out within the variable range of the PSI, focusing on the region where its effects are relatively pronounced [21]. As shown in the figure, for the MaxCut3 problem composed solely of anti-Ferro-type interactions, both the behavior of the pseudo-spins (Figs. 6(a), 6(c)) and the behavior of the Ising energy (IE) (Figs. 6(b), 6(d)) were confirmed to exhibit the expected transition behavior to the stable state on the IE, consistent with our predictions. On the other hand, for the complete graph problem () where Ferro and Anti-Ferro interactions coexist, it was confirmed that when PSI exceeds a certain threshold (roughly ), the pseudo-spins are drawn toward the 0-state (Fig. 7(c)), and at , all pseudo-spins ultimately end up in the 0-state (Fig. 7(e)), and and regarding behavior on the Ising energy, while it temporarily stabilizes, it later shifts toward an unstable state and converges to a specific energy level (Fig. 7(d), (f)), exhibiting unexpected behavior.
@@
Figure 8 graphically illustrates the pseudo-spin () dependence of PSI when . The PSI is defined on the basis of the optical interference phenomenon and is described by the amplitudes of the two interacting pseudo-spins. Consequently, the pseudo-spin dependence of PSI exhibits the nonlinearity and asymmetry shown in the figure. The blue and red lines on the graph indicate the positions of combinations where PSI is 0. Roughly speaking, the effect of PSI on acts to move from any combination state of toward the direction of its nearest position of a combination where PSI is 0 (i.e. toward the direction of these blue and red lines). For Anti-Ferro-type PSI, the PSI acting on will ultimately be more effective for transitions toward = (0, 1) than for transitions toward = (1, 0). This arises because, in the presence of the competing PSI, the asymmetry of the PSI allows the vicinity of = (0, 1) to approach the Blue-line more rapidly than the vicinity of = (1, 0) (see the Anti-Ferro-type PSI shown in Fig.8(a) and 8(b)). Consequently, the dynamics-driven attraction becomes more effective. However, focusing on the transition effect that acts on , the opposite holds true, so this asymmetry cancels out and its effectiveness disappears. This is considered the reason why the computational simulation results in the MaxCut3 problem exhibit the behavior we initially expected (see Fig.6). On the other hand, in the case of Ferro-type PSI, the transition effects acting on both and are more effective overall when they transition toward = (0, 0) is more effective than the transition effect toward = (1, 1) (see Ferro-type PSI shown in Figs. 8(c) and 8(d)), and unlike anti-Ferro-type PSI, there is no mutual cancelation, so the dynamics-driven attraction toward the 0-state becomes more effective. This can be considered a major factor in the phenomenon in which all pseudo-spins in with Ferro-type PSI transition to the 0-state (see Fig.7(e)). Furthermore, it is thought that the balance between the dynamics-driven attraction toward the 0-state&1-state and the magnitude of the PSI effect leads to more complex behavior and consequently different results (see Figs.7(a),(c),(e)). This result predicted by numerical simulations — namely, that in a many-body-like system created by symmetry breaking, dual elements expected to be nearly balanced in population instead exhibit behavior where only one element is created — is not only reminiscent of the physical process called baryogenesis (an unsolved theoretical research subject in physical cosmology [24]), but is also particularly intriguing as it represents an observable singular case. However, it is clear that such effects become a major obstacle when considering the application of SSBM to combinatorial optimization problems. Moreover, if the above predictions and hypotheses are correct, then a solution can be proposed relatively easily. Specifically, it is thought that this asymmetry can be resolved by constructing a system conjugate to the one examined thus far and averaging it with the original system. The crucial point here is that both this conjugate system and the system averaged together with the conjugate system are physically implementable, just like the conventional SSBM. Here, the equations governing this composite system can be described as follows.
@@
Under the same conditions (see Figs.6 and 7), the verification of numerical simulation based on Eq.(8) revealed that while the case applied to MaxCut3 showed almost no difference from the conventional method, the case applied to exhibited a drastic difference. Figure 9 shows the results of the numerical simulations concerning . The phenomenon of avalanche-like transition to the 0-state observed in Fig.7(c)&(e) has disappeared, and stabilization, as measured by the IE, was shown to proceed without hindering. These results clearly demonstrate that the aforementioned predictions and hypotheses are correct. Furthermore, in the following subsection, we will continue our discussion based on the composite system described by Eq.(8).

### v2

@@ Abstract
In a previous paper, we proposed a unique physically implemented type simulator for combinatorial optimization problems, called a spontaneous symmetry breaking machine (SSBM). In this paper, we first report the results of the experimental verification of SSBM using a small-scale benchmark system, and then describe numerical simulations using the benchmark problems () conducted to confirm its usefulness for large-scale problems. From 1000 samples with different initial fluctuations, it became clear that SSBM can explore a single extremely stable state. This unique behavior holds the potential to become a significant advantage over other simulators.
@@ I Introduction
Symmetry and its (spontaneous) breaking are one of the most significant concepts for understanding various natural phenomena, and it is widely known that they play an essential role in a wide range of areas covered by science [1]. In applied physics, which is at the heart of scientific applications, and in condensed matter physics, which is the main foundation of applied physics, there are many arguments suggesting that spontaneous symmetry breaking (SSB) may play an important role in various phenomena. In those, the discussion is mainly focused on explaining phase transitions (e.g. spontaneous magnetization [2], insulator-metal phases [3]) and changes of state before and after phase transitions, and the main emphasis is on understanding the manifestation of macroscopic properties based on the assumption of the existence of microscopic elements (e.g. electron spin). In dissipative systems, concepts such as dissipative structure and synergetics have been discovered, leading to a broader and deeper understanding of natural phenomena from the perspective of dynamical macroscopic orders created by symmetry breaking [4, 5]. In the field of elementary particle theory, it goes even further and plays an important role in understanding the manifestation of elementary particles (microscopic elements) themselves such as the Higgs boson [6, 7].
In recent years, there has been growing interest in attempts to solve problems in an extremely wide range of fields (logistics, financial portfolios, drug discovery, machine learning, etc.) by treating them as combinatorial optimization problems (COPs). At the same time, there has been a lot of research focused on Ising problems that can be mapped to COPs in the NP-hard category [8]. In particular, quantum annealers (a type of quantum computer) [9, 10] and coherent Ising machines (based on laser oscillation phenomena) [11] that use physically implemented type simulators to explore the ground state of Ising model systems are attracting attention.
@@ II SSB machine
Polarized coherent clock pulses with pulse width and repetition interval are input to the optical input port of a 12 MZM used as a partial inflow gate. The -th output light pulse from output port A of the MZM is converted into an electrical pulse by a photodiode after passing through an optical variable delay line to adjust the delay and timing. This electrical pulse is then expanded to a pulse width of using a Bessel filter, while its power is compensated by two amplifiers while suppressing excessive noise, and is used for intensity modulation by the MZM for the ()-th optical input pulse. In the above operation of this circuit, we can identify independent fully dissipative systems defined by virtual boundaries where the -th and ()-th clock pulses interact, with the MZM serving as the common inflow gate. And, when the condition is satisfied, this fully dissipative system maintains an extremely singular state described by the following iterative equation using the parameters phase difference-induced efficiency , the transmittance and to port A at the input of the clock pulses and ()-th, and the static phase condition (offset) , related to MZM, thus establishing the dissipative causality.
Here, we set the remaining degrees of freedom as and . These are the parameters that enable the SSB phenomenon to occur in the SSBM [12, 22]. (Unlike examples that realize systems following similar iterative equations by incorporating analog functions into the digital processing flow using MZM characteristics [23], robust causality as a phenomenon is realized here.) However, when the optical clock pulse is supplied and the dissipative causality is established, the states appear and are maintained (becoming hidden states). Therefore, it is necessary to input an initialization optical pulse train from the optical input port with adjusted timing to transition all full-dissipative systems to an unstable fixed point (), which is the state before SSB begins (SBSB). Moreover, we will consider the construction of a many-body-like system by imparting interrelationships between full-dissipative systems capable of causing the SSB phenomenon described above. Specifically, we treat the two attractors in the above full-dissipative systems as pseudo spin up and spin down, and introduce pseudo-spin interactions (pSIs) represented by the following equations, which can be implemented by an optical interference circuit.
@@
As is clear from the equations these pSIs are different from the so-called exchange interactions, but they satisfy the behavior consistent with the interpretation of exchange energy (namely, ferromagnetic types stabilize when states are aligned, while antiferromagnetic types stabilize when states are opposite). This is necessary to deal with the fact that pseudo-spins are continuous variables rather than discrete variables. Even though pseudo-spins are continuous variables, if we use interactions that follow exchange interactions as they are, we cannot guaranty the expected behavior in regions other than those corresponding to the values assigned to spins. The introduction of these pSIs yields new and interesting results. As is clear from Equations (4) and (5), when full-dissipative systems are simultaneously in the SBSB , the effects of pSIs disappear. Therefore, it is possible to transition the entire complex system consisting of full-dissipative systems with pSIs to SBSB through the initialization process described above. Moreover, this complex system creates a many-body-like system consisting of pseudo-spins, incorporating the effects of pSIs through the SSB phenomenon. We focus on the phenomenon where the symmetry temporarily restored by the above initialization process spontaneously breaks and transitions to a stable fixed point (i.e., the SSB phenomenon), utilizing this as a mechanics-derived attraction driving the transition between two stable fixed points. (Despite the circuit configuration similarity to coherent Ising machines using MZM in the feedback loop [24], it is important to note that SSBM employs attraction derived from dynamics—rather than utilizing oscillatory phenomena—as the crucial driving principle that leads to either the up or down state.) We refer to a many-body-like system composed of pseudo-spins as described above as an Ising-model-like system.
@@ III Proof-of-principle experiment
To understand the creation of the Ising-model-like system through the SSB phenomenon, we cannot simply apply the concepts developed in condensed matter physics. And we do not know of any other specific examples of the phenomenon of many-body-like systems being created by SSB. However, since consistency with the interpretation of exchange interaction energy was ensured when introducing pSIs, it is expected that a correspondence will exist between the stable states in the Ising model system and those in the Ising-model-like system. Nevertheless, the correspondence with the Ising model system discussed [25, 26]. regarding continuous-variable Ising machines utilizing bifurcation phenomena through feedback strength control cannot be directly applied. In our previous paper, we conducted numerical simulations to confirm this correspondence and found it promising [12]. Here, we present one of the most compelling pieces of evidence to ensure the reliability of this correspondence (duality) by experimentally demonstrating the effectiveness of SSBM as a simulator for COPs.
@@
Figure 3 is a third-order graph representing a COP belonging to the category called MaxCut3 [15], which is addressed in the proof-of-principle experiment. This COP is an established problem that was used in demonstration experiments of coherent Ising machines, which are one type of physically implemented Ising machine. And this COP has the advantage that dynamic switching is not required in the optical delay interference circuit (ODIC) [27, 28] for the physical implementation of pseudo-spin interactions.
@@
Figure 4(a) shows a schematic diagram of the ODIC, and Fig.4(b) shows a schematic diagram of a proof-of-principle system for SSBM that implements this circuit. Here, the ODIC was fabricated as a silica-based planar lightwave circuit, while the other parts of the system were constructed using commercially available components. The ODIC has input ports corresponding to the output port and the complementary output port of the 12 MZM, and the optical delay interference between the input optical pulses from these input ports physically implements the relationships (pseudo-spin interactions) between the 16 full-dissipative systems, which is then output to the variable optical delay line in the subsequent stage. This ODIC is designed so that, under the condition that no additional optical amplification is performed, the losses for branching and adjustment (phase, attenuation) are taken into account to satisfy . The delay of the third route corresponding to is set to be the smallest, and the delays of the routes corresponding to the reference route (main route and second route:), and are set to be times larger than the clock pulse interval. In addition, by adjusting the phase shifters and attenuators, the pseudo-spin interaction represented by Eqns.(2)-(7) is physically implemented through optical delay interference. A series of optical clock pulses with a repetition frequency of 1.02 GHz is extracted from a series of optical output pulses from an active mode-locked laser with a repetition frequency of 8.16 GHz and a pulse width of 10 ps and then split into two. To allow repeated trials, optical coherent clock pulses with a width of and a period of 0.996 MHz are extracted from one of these optical pulse trains and reset regions with a width of are set. From the other of these optical pulse trains, a sequence consisting of the first 16 consecutive optical pulses is extracted and prepared as an optical initializing pulse train. In addition, we adjusted the timing so that the electrical pulse originating from the first optical pulse in this optical initializing pulse train is input as the intensity-modulated signal to the 12 MZM at the same time as the first optical pulse of the optical coherent clock pulses is input to the 12 MZM. With the above settings, SSBM will reset and immediately transition to the NS, causing the SSB phenomenon to start, repeating this behavior. The peak power of the inflow optical pulse detected by the photodiode corresponds to the transmittance of each system, so this electrical pulse was normalized to monitor the behavior of the system. Using a real-time oscilloscope, a series of 800 trials was recorded at once.
@@
The waveforms shown in Fig. 5(a) correspond to the case where the pSI is turned off (set ) using attenuators in the ODIC. It was confirmed that the fundamental SSB phenomenon occurs in each of the 16 independent full-dissipative systems, where the system transitions to either the 0-state or 1-state in a manner that reflects the initial fluctuations. In cases without pSI, since full-dissipative systems are entirely independent and do not compete with each other, state transitions are completed in approximately 20 steps, with little variation in the number of steps required. Figures 5(b)-(c) show the normalized monitoring signals of the SSBM with pseudo-spin interactions. In this example, it can be confirmed that the state transition is completed in approximately 30 steps, successfully transitioning to the state corresponding to the most stable solution. Furthermore, the number of steps required for solution search exhibited a relatively large variation and required significantly more steps compared to when pSI was off, with some cases requiring over 50 steps. This variability is believed to stem from the fact that pSI-induced conflicts exhibit different behaviors in individual cases. Based on data obtained from experiments, when thresholding was performed defining values above 80% of the normalized value as up and values below 20% as down, the success rate to find the most stable solution was approximately 97%. These results demonstrate that SSBM is a promising physical implementation simulator for combinatorial optimization problems. They also strongly support the existence of a duality (correspondence) between the Ising-model-like system states that appear in SSBM and the stable states of the Ising model system.
@@ IV.1 Asymmetric Effects in Pseudo-Spin Interactions
Figures 6 and 7 show the results of numerical simulations for the MaxCut3 problem and a fully connected complete graph problem, respectively. The significant difference in the magnitude of the pSI employed for these two problems is due to the differing densities of coupling, and the calculations were carried out within the variable range of the pSI, focusing on the region where its effects are relatively pronounced [29]. As shown in the figure, for the MaxCut3 problem composed solely of anti-Ferro-type interactions, both the behavior of the pseudo-spins (Figs. 6(a), 6(c)) and the behavior of the Ising energy (Figs. 6(b), 6(d)) were confirmed to exhibit the expected transition behavior to the stable state on the Ising energy, consistent with our predictions. On the other hand, for the complete graph problem () where Ferro and Anti-Ferro interactions coexist, it was confirmed that when pSI exceeds a certain threshold (roughly ), the pseudo-spins are drawn toward the 0-state (Fig. 7(c)), and at , all pseudo-spins ultimately end up in the 0-state (Fig. 7(e)), and regarding behavior on the Ising energy, while it temporarily stabilizes, it later shifts toward an unstable state and converges to a specific energy level (Fig. 7(d), (f)), exhibiting unexpected behavior.
@@
Figure 8 graphically illustrates the pseudo-spin () dependence of pSI when . The pSI is defined on the basis of the optical interference phenomenon and is described by the amplitudes of the two interacting pseudo-spins. Consequently, the pseudo-spin dependence of pSI exhibits the nonlinearity and asymmetry shown in the Fig.8. The blue and red lines on the graph indicate the positions of combinations where pSI is 0. Roughly speaking, the effect of pSI on acts to move from any combination state of toward the direction of its nearest position of a combination where pSI is 0 (i.e. toward the direction of these blue and red lines). For Anti-Ferro-type pSI, the pSI acting on will ultimately be more effective for transitions toward = (0, 1) than for transitions toward = (1, 0). This arises because, in the presence of the competing pSI, the asymmetry of the pSI allows the vicinity of = (0, 1) to approach the Blue-line more rapidly than the vicinity of = (1, 0) (see the Anti-Ferro-type pSI shown in Fig.8(a) and 8(b)). Consequently, the dynamics-derived attraction becomes more effective. However, focusing on the transition effect that acts on , the opposite holds true, so this asymmetry cancels out and its effectiveness disappears. This is considered the reason why the computational simulation results in the MaxCut3 problem exhibit the behavior we initially expected (see Fig.6). On the other hand, in the case of Ferro-type pSI, the transition effects acting on both and are more effective overall when they transition toward = (0, 0) is more effective than the transition effect toward = (1, 1) (see Ferro-type pSI shown in Figs. 8(c) and 8(d)), and unlike anti-Ferro-type pSI, there is no mutual cancelation, so the dynamics-derived attraction toward the 0-state becomes more effective. This can be considered a major factor in the phenomenon in which all pseudo-spins in with Ferro-type pSI transition to the 0-state (see Fig.7(e)). Furthermore, it is thought that the balance between the dynamics-derived attraction toward the 0-state&1-state and the magnitude of the pSI effect leads to more complex behavior and consequently different results (see Figs.7(a),(c),(e)). This unexpected phenomenon, while highly intriguing [30], it is clear that such effects become a major obstacle when considering the application of SSBM to combinatorial optimization problems. Moreover, if the above predictions and hypotheses are correct, then a solution can be proposed relatively easily. Specifically, it is thought that this asymmetry can be resolved by constructing a system conjugate to the one examined thus far and averaging it with the original system. The crucial point here is that both this conjugate system and the system averaged together with the conjugate system are physically implementable, just like the conventional SSBM. Here, the equations governing this composite system can be described as follows.
@@
Under the same conditions (see Figs.6 and 7), the verification of numerical simulation based on Eq.(8) revealed that while the case applied to MaxCut3 showed almost no difference from the conventional method, the case applied to exhibited a drastic difference. Figure 9 shows the results of the numerical simulations concerning . The phenomenon of avalanche-like transition to the 0-state observed in Fig.7(c)&(e) has disappeared, and stabilization, as measured by the Ising energy, was shown to proceed without hindering. These results clearly demonstrate that the aforementioned predictions and hypotheses are correct. Furthermore, in the following subsection, we will continue our discussion based on the composite system described by Eq.(8).